Indicate the oxidation step of the central atom for the ionic complex compound [CrCl2(NH3)4]+
I know the answer is +3, but why?
1
Expert's answer
2018-05-18T02:51:17-0400
Central atom - Cr. The oxidation state of Cl is -1, the oxidation state of NH3 is 0, because it is neutral ligand, total charge of ionic complex is +1. We can find the oxidation state of Cr by solving equation:
x + (-1)*2 + 0*4 = +1
x - 2 = 1
x = 1+2 = 3